# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ C.2.1 Combination of random variables
$\sqrt{\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)^{2}}\right)-\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)}\right)^{2}}=\mathcal{O}\left(\frac{1}{\sqrt{N_{2}}}\right)\,.$ (49)
By this we conclude that, in the $N_{2}\rightarrow\infty$ limit, the distribution of the r.v. $\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)$ narrows around the mean value $\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)}\right)$.
We can proceed analogously to evaluate the mean value fluctuations; therefore we have to repeat the same analysis on the mean of $p_{h^{(3)}_{i}}^{(\chi)}\left({x}\right)$, i.e. (from Eq. (43)) $\sum_{j=1}^{N_{2}}w^{(2)}_{ij}\left\langle{g^{(2)}_{j}}\right\rangle$. In this case, we have
$\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\sum_{j=1}^{N_{2}}w^{(2)}_{ij}\left\langle{g^{(2)}_{j}}\right\rangle}\right)=\sum_{j=1}^{N_{2}}\underbrace{\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({w^{(2)}_{ij}}\right)}_{=0}\left\langle{g^{(2)}_{j}}\right\rangle=0\,.$ (50)
@@
$\sqrt{\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{h^{(3)}_{i}}\right\rangle^{2}}\right)-\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{h^{(3)}_{i}}\right\rangle}\right)^{2}}=\mathcal{O}(1)\,.$ (51)
This means that the center of the distribution $p_{h^{(3)}_{i}}^{\chi}\left({x}\right)$ keeps fluctuating from node to node even in the limit $N_{2}\rightarrow\infty$.
@@
$\left(\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{h^{(3)}_{i}}\right\rangle^{2}}\right)-\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{h^{(3)}_{i}}\right\rangle}\right)^{2}\right)=\sigma^{2}_{w}\frac{1}{N_{2}}\sum_{j=1}^{N_{2}}\left\langle{g^{(2)}_{j}}\right\rangle^{2}\xrightarrow{N_{2}\rightarrow\infty}\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\left\langle{g^{(2)}_{j}}\right\rangle^{2}}\right)\,.$ (52)
Also, it is easy to show that the set $\left\{w^{(2)}_{ij}\left\langle{g^{(2)}_{j}}\right\rangle\right\}_{j=1}^{N_{2}}$ respect the conditions of Thm. C.3. Therefore, to summarize
$p_{\left\langle{h^{(3)}_{i}}\right\rangle}\left({x}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(x;0,\sigma^{2}_{w}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\left\langle{g^{(2)}_{j}}\right\rangle^{2}}\right)\right)\,$, (53)
@@
$p_{\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)}\left({x}\right)\xrightarrow{N_{2}\rightarrow\infty}\delta\left(x-\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)}\right)\right)\,$. (54)
@@ Appendix D Derivation of the distributions
In this section, our focus is on the r.v. $O^{(c)}$. The primary goal is to prove Thm. 4.1 by deriving expressions for $p_{\mu_{c}}\left({x}\right)$ and $p_{O^{(c)}}^{(\chi)}\left({x}\right)$. These expressions are crucial for determining $p_{f_{0}}\left({x}\right)$, as explored in App. E.1.
@@
  • In App. D.2, we will derive an expression for $p_{\mu_{c}}\left({x}\right)$ for various cases of interest.
  • In App. D.3, we will derive $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ for the same cases.
@@
Consider the r.v. $h^{(1)}_{i}$ defined as:
$h^{(1)}_{i}=\sum_{j}w^{(0)}_{ij}\xi_{j}.$ (55)
The independence of $h^{(1)}_{i}$ is a direct consequence of the assumed independence of the input components $\xi_{j}$.
For the r.v. $O^{(c)}$, defined by
$O^{(c)}\left(\boldsymbol{\xi};\mathcal{W}\right)=\sum_{m=1}^{N_{1}}w^{(1)}_{cm}g^{(1)}_{m}$, (56)
we require independence in the set $\{g^{(1)}_{m}\}$ to apply the Central Limit Theorem (CLT). Since the activation function is an element-wise operation on $\{h^{(1)}_{m}\}$, the independence of $h^{(1)}_{m}$ suffices.
Jointly normally distributed random variables are independent if and only if their covariance is zero. We calculate the covariance $\textrm{Cov}(h^{(1)}_{i},h^{(1)}_{j})$ and demonstrate that it converges to 0 w.h.p. as the input dimensionality $d\rightarrow\infty$. Specifically, we have:
@@
Fixing the set $\mathcal{W}$, $h^{(1)}_{i}$ is a linear combination of r.v.s which meet the conditions of Thm. C.3. Thus, in the limit as $d\rightarrow\infty$, we have
$p_{h^{(1)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}\left(x;\sum_{j}w^{(0)}_{ij}\left\langle{\xi_{j}}\right\rangle,\sum_{j}\left(w^{(0)}_{ij}\right)^{2}\sigma^{2}_{\xi_{j}}\right)=\mathcal{N}\left(x;0,\sum_{j}\left(w^{(0)}_{ij}\right)^{2}\right),$ (62)
@@
$\sum_{j=1}^{d}\left(w^{(0)}_{ij}\right)^{2}=d\;\mkern 1.5mu\overline{\mkern-1.5mu\left(w^{(0)}_{ij}\right)^{2}\mkern-1.5mu}\mkern 1.5mu\xrightarrow{d\rightarrow\infty}\sigma^{2}_{w}\Longrightarrow p_{h^{(1)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}\left(x;0,\sigma^{2}_{w}\right).$ (63)
Now, passing through the activation function, and possibly through the pooling layer, we have a set of i.i.d. variables, $\{\rho^{(1;m)}_{l}\}_{l=1}^{\lceil\frac{N_{1}}{m}\rceil}$. Furthermore, by hypothesis, the set $\{\rho^{(1;m)}_{l}\}_{l=1}^{\lceil\frac{N_{1}}{m}\rceil}$ satisfies the conditions of Thm. C.3. Therefore, the distribution of the r.v.
$O^{(c)}\left(\boldsymbol{\xi};\mathcal{W}\right)=\sum_{l=1}^{N_{1}}w^{(1)}_{cl}\rho^{(1;m)}_{l}$. (64)
@@
$\lim_{N_{1}\rightarrow\infty}p_{O^{(c)}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\mu_{c},\textrm{Var}_{\chi}\left({O^{(c)}}\right)\right).$ (65)
In particular, since the variables $\{\rho^{(1;m)}_{l}\}_{l=1}^{\lceil\frac{N_{1}}{m}\rceil}$ are i.i.d. and the weights $\{w^{(1)}_{cl}\}_{l=1}^{\lceil\frac{N_{1}}{m}\rceil}$ are fixed constants from initialization, defining
$\lim_{D\to\infty}\left\langle{\rho^{(1;m)}_{l}}\right\rangle=\left\langle{\rho^{(1;m)}}\right\rangle\;\;\;\forall l,$ (66)
$\lim_{D\to\infty}\textrm{Var}_{\chi}\left({\rho^{(1;m)}_{l}}\right)=\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)\;\;\;\forall l.$ (67)
@@
$\mu_{c}$ $=\left\langle{\rho^{(1;m)}}\right\rangle\sum_{l=1}^{N_{1}}w^{(1)}_{cl}\equiv\left\langle{\rho^{(1;m)}}\right\rangle S_{w_{c}}^{(N_{1})}$ (68)
$\textrm{Var}_{\chi}\left({O^{(c)}}\right)$ $=\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)\sum_{l=1}^{N_{1}}\left(w^{(1)}_{cl}\right)^{2}\equiv\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)S_{w_{c}^{2}}^{(N_{1})}.$ (69)
@@
$\lim_{N_{1}\rightarrow\infty}\textrm{Var}_{\chi}\left({O^{(c)}}\right)=\sigma_{w}^{2}\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)\;\;\forall c.$ (70)
@@
  • $\left\langle{\rho^{(1;m)}}\right\rangle=0$:
@@
    $\lim_{N_{1}\rightarrow\infty}p_{O^{(c)}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;0,\sigma_{w}^{2}\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)\right).$ (71)
  • $\left\langle{\rho^{(1;m)}}\right\rangle\neq 0$:
@@
    $\lim_{N_{1}\rightarrow\infty}p_{O^{(c)}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\left\langle{\rho^{(1;m)}}\right\rangle S_{w_{c}}^{(N_{1})},\sigma_{w}^{2}\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)\right).$ (72)
@@
    $p_{\mu_{c}}\left({x}\right)=\mathcal{N}\left(x;0,\sigma_{w}^{2}\left\langle{\rho^{(1;m)}}\right\rangle^{2}\right).$ (74)
D.2 Derivation of $\boldsymbol{p_{\mu_{c}}\left({x}\right)}$ for some emblematic cases
Thm. D.1 demonstrates that the decisive element for deriving $p_{\mu_{c}}\left({x}\right)$ is $\left\langle{\rho^{(1;m)}}\right\rangle$. To make the discussion more concrete, we will consider some cases to illustrate their distinguishing features. In particular, we will use the same settings discussed at the beginning of Sec. 5.
@@ D.2 Derivation of $\boldsymbol{p_{\mu_{c}}\left({x}\right)}$ for some emblematic cases
$\rho^{(1;1)}_{i}=\rho^{(1;1)}\left({g\left({h^{(1)}_{i}}\right)}\right)=h^{(1)}_{i}$. (75)
From Eq.s (62) and (75), it follows that $\left\langle{\rho^{(1;1)}}\right\rangle=0$. Therefore, in this setting, we have no presence of IGB, and according to Eq. 60, we have:
$\lim_{D\rightarrow\infty}p_{\mu_{c}}\left({x}\right)=\delta(x).$ (76)
Note that (76) is formally true only in the limit of an infinite size dataset. For a finite size dataset, we must compute the average $\left\langle{\rho^{(1;1)}}\right\rangle$ using the empirical distribution, $p_{h^{(1)}_{k}}^{(\chi;E)}\left({x}\right)\equiv\sum_{a=1}^{D}\frac{1}{D}\delta\left(h^{(1)}_{k}-\hat{h}_{k}^{(1)}\left(\boldsymbol{\xi}^{(a)}\right)\right)$, where the set $\left\{\hat{h}_{k}^{(1)}\left(\boldsymbol{\xi}^{(a)}\right)\right\}_{a=1}^{D}$ contains the values mapped from each element of the dataset to the $k^{\text{th}}$ node of the first hidden layer. We then have
$\begin{split}&\left\langle{\rho^{(1;1)}_{j}}\right\rangle=\int_{\mathbb{R}}x\,p_{h^{(1)}_{k}}^{(\chi;E)}\left({x}\right)dx=\frac{1}{D}\sum_{a=1}^{D}\hat{h}_{j}^{(1)}\left(\boldsymbol{\xi}^{(a)}\right)\underset{\boldsymbol{A}}{\simeq}\frac{1}{\sqrt{D}}Z_{j}=\tilde{Z}_{j},\end{split}$ (77)
@@
Therefore, considering finite size effects, $\left\langle{\rho^{(1;1)}}\right\rangle\neq\left\langle{\rho^{(1;1)}_{j}}\right\rangle$. In this case, instead of starting from Eq. (68), we can proceed similarly to the derivation of Eq. (LABEL:eq:<h>^l+1), obtaining
$p_{\mu_{c}}\left({x}\right)=\mathcal{N}\left(x;0,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\tilde{Z}_{j}^{2}\mkern-1.5mu}\mkern 1.5mu\right)=\mathcal{N}\left(x;0,\frac{4}{D}\right).$ (78)
Note that in step $\boldsymbol{A}$ of Eq. (77), we again used the CLT for the sum of r.v.s. The key difference with the previous derivation is that we do not use the CLT for $p_{h^{(1)}_{k}}^{(\chi)}\left({x}\right)$; instead, we substitute the empirical distribution, $p_{h^{(1)}_{k}}^{(\chi;E)}\left({x}\right)$. This latter distribution, defined on the dataset elements, accounts for the finite size effects of the dataset itself (see dependence from $D$). Note that the distribution of the variables involved in the sum in the last part of Eq. (77) ($\tilde{Z}_{j}$) narrows as $D\rightarrow\infty$; in this limit, the finite size effects disappear, and we converge to the result in Eq. (76).
@@
$\rho^{(1;1)}_{i}=\rho^{(1;1)}\left({g^{(1)}_{j}}\right)=g^{(1)}_{j}=g\left(h^{(1)}_{j}\right)\equiv\max\{0,h^{(1)}_{j}\}$. (79)
$g^{(1)}_{j}$ will follow the same distribution as $h^{(1)}_{j}$ on the positive support, since $g^{(1)}_{j}=h^{(1)}_{j}$ for $h^{(1)}_{j}>0$. The probability density on the negative support of $p_{h^{(1)}_{j}}^{(\chi)}\left({x}\right)$ will collapse to 0 since $g^{(1)}_{j}=0$ for $h^{(1)}_{j}<0$. Substituting $p_{h^{(1)}_{j}}^{(\chi)}\left({x}\right)$ with the Gaussian distribution from Eq. (63), we get:
$\begin{split}&p_{g^{(1)}_{j}}^{(\chi)}\left({x}\right)=\left(\int_{\mathbb{R}^{-}}p_{h^{(1)}_{j}}^{(\chi)}\left({\tilde{x}}\right)d\tilde{x}\right)\delta\left(x\right)+\Theta\left(x\right)p_{h^{(1)}_{j}}^{(\chi)}\left({x}\right)=\frac{1}{2}\delta\left(x\right)+\Theta\left(x\right)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right),\end{split}$ (80)

### v5

@@ C.2.1 Combination of random variables
$\sqrt{\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)^{2}}\right)-\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)}\right)^{2}}=\mathcal{O}\left(\frac{1}{\sqrt{N_{2}}}\right)\,.$ (49)
By this we conclude that, in the $N_{2}\rightarrow\infty$ limit, the distribution of the r.v. $\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)$ narrows around the mean value $\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)}\right)$.
We can proceed analogously to evaluate the mean value fluctuations; therefore we have to repeat the same analysis on the mean of $f_{Z^{(3)}_{i}}^{(\chi)}\left({z{}{}}\right)$, i.e. (from Eq. (43)) $\sum_{j=1}^{N_{2}}w^{(2)}_{ij}\left\langle{H^{(2)}_{j}}\right\rangle$. In this case, we have
$\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\sum_{j=1}^{N_{2}}w^{(2)}_{ij}\left\langle{H^{(2)}_{j}}\right\rangle}\right)=\sum_{j=1}^{N_{2}}\underbrace{\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({w^{(2)}_{ij}}\right)}_{=0}\left\langle{H^{(2)}_{j}}\right\rangle=0\,.$ (50)
@@
$\sqrt{\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{Z^{(3)}_{i}}\right\rangle^{2}}\right)-\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{Z^{(3)}_{i}}\right\rangle}\right)^{2}}=\mathcal{O}(1)\,.$ (51)
This means that the center of the distribution $f_{Z^{(3)}_{i}}^{(\chi)}\left({z{}{}}\right)$ keeps fluctuating from node to node even in the limit $N_{2}\rightarrow\infty$.
@@
$\left(\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{Z^{(3)}_{i}}\right\rangle^{2}}\right)-\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left\langle{Z^{(3)}_{i}}\right\rangle}\right)^{2}\right)=\sigma^{2}_{w}\frac{1}{N_{2}}\sum_{j=1}^{N_{2}}\left\langle{H^{(2)}_{j}}\right\rangle^{2}\xrightarrow{N_{2}\rightarrow\infty}\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\left\langle{H^{(2)}_{j}}\right\rangle^{2}}\right)\,.$ (52)
Also, it is easy to show that the set $\left\{w^{(2)}_{ij}\left\langle{H^{(2)}_{j}}\right\rangle\right\}_{j=1}^{N_{2}}$ respect the conditions of Thm. C.3. Therefore, to summarize,
$f_{\left\langle{Z^{(3)}_{i}}\right\rangle}\left({m_{z}}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(m_{z};0,\sigma^{2}_{w}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\left\langle{H^{(2)}_{j}}\right\rangle^{2}}\right)\right)\,$, (53)
@@
$f_{\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)}\left({v_{z}}\right)\xrightarrow{N_{2}\rightarrow\infty}\delta\left(v_{z}-\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)}\right)\right)\,$. (54)
@@ Appendix D Derivation of the distributions
In this section, our focus is on the r.v. $O_{c}$. The primary goal is to prove Thm. 4.1 by deriving expressions for $f_{\mu_{c}}\left({m}\right)$ and $f_{O_{c}}^{(\chi)}\left({o}\right)$. These expressions are crucial for determining $f_{G_{0}}\left({g_{0}}\right)$, as explored in App. E.1.
@@
  • In App. D.2, we will derive an expression for $f_{\mu_{c}}\left({m}\right)$ for various cases of interest.
  • In App. D.3, we will derive $f_{O_{c}}^{(\chi)}\left({o}\right)$ for the same cases.
@@
Consider the r.v. $Z^{(1)}_{i}$ defined as:
$Z^{(1)}_{i}=\sum_{j}w^{(0)}_{ij}\xi_{j}.$ (55)
The independence of $Z^{(1)}_{i}$ is a direct consequence of the assumed independence of the input components $\xi_{j}$.
For the r.v. $O_{c}$, defined by
$O_{c}\left(\boldsymbol{\xi};\mathcal{W}\right)=\sum_{m=1}^{N_{1}}w^{(1)}_{cm}H^{(1)}_{m}$, (56)
we require independence in the set $\{H^{(1)}_{m}\}$ to apply the Central Limit Theorem (CLT). Since the activation function is an element-wise operation on $\{Z^{(1)}_{m}\}$, the independence of $Z^{(1)}_{m}$ suffices.
Jointly normally distributed random variables are independent if and only if their covariance is zero. We calculate the covariance $\textrm{Cov}(Z^{(1)}_{i},Z^{(1)}_{j})$ and demonstrate that it converges to 0 w.h.p. as the input dimensionality $d\rightarrow\infty$. Specifically, we have:
@@
Fixing the set $\mathcal{W}$, $Z^{(1)}_{i}$ is a linear combination of r.v.s which meet the conditions of Thm. C.3. Thus, in the limit as $d\rightarrow\infty$, we have
$f_{Z^{(1)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}\left(z{}{};\sum_{j}w^{(0)}_{ij}\left\langle{\xi_{j}}\right\rangle,\sum_{j}\left(w^{(0)}_{ij}\right)^{2}\sigma^{2}_{\xi_{j}}\right)=\mathcal{N}\left(z{}{};0,\sum_{j}\left(w^{(0)}_{ij}\right)^{2}\right),$ (62)
@@
$\sum_{j=1}^{d}\left(w^{(0)}_{ij}\right)^{2}=d\;\mkern 1.5mu\overline{\mkern-1.5mu\left(w^{(0)}_{ij}\right)^{2}\mkern-1.5mu}\mkern 1.5mu\xrightarrow{d\rightarrow\infty}\sigma^{2}_{w}\Longrightarrow f_{Z^{(1)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}\left(z{}{};0,\sigma^{2}_{w}\right).$ (63)
Now, passing through the activation function, and possibly through the pooling layer, we have a set of i.i.d. variables, $\{U^{(1;k)}_{l}\}_{l=1}^{\lceil\frac{N_{1}}{k}\rceil}$. Furthermore, by hypothesis, the set $\{U^{(1;k)}_{l}\}_{l=1}^{\lceil\frac{N_{1}}{k}\rceil}$ satisfies the conditions of Thm. C.3. Therefore, the distribution of the r.v.
$O_{c}\left(\boldsymbol{\xi};\mathcal{W}\right)=\sum_{l=1}^{N_{1}}w^{(1)}_{cl}U^{(1;k)}_{l}$. (64)
@@
$\lim_{N_{1}\rightarrow\infty}f_{O_{c}}^{(\chi)}\left({o}\right)=\mathcal{N}\left(o;\mu_{c},\textrm{Var}_{\chi}\left({O_{c}}\right)\right).$ (65)
In particular, since the variables $\{U^{(1;k)}_{l}\}_{l=1}^{\lceil\frac{N_{1}}{k}\rceil}$ are i.i.d. and the weights $\{w^{(1)}_{cl}\}_{l=1}^{\lceil\frac{N_{1}}{k}\rceil}$ are fixed constants from initialization, defining
$\lim_{D\to\infty}\left\langle{U^{(1;k)}_{l}}\right\rangle=\left\langle{U^{(1;k)}}\right\rangle\;\;\;\forall l,$ (66)
$\lim_{D\to\infty}\textrm{Var}_{\chi}\left({U^{(1;k)}_{l}}\right)=\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)\;\;\;\forall l.$ (67)
@@
$\mu_{c}$ $=\left\langle{U^{(1;k)}}\right\rangle\sum_{l=1}^{N_{1}}w^{(1)}_{cl}\equiv\left\langle{U^{(1;k)}}\right\rangle S_{w_{c}}^{(N_{1})}$ (68)
$\textrm{Var}_{\chi}\left({O_{c}}\right)$ $=\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)\sum_{l=1}^{N_{1}}\left(w^{(1)}_{cl}\right)^{2}\equiv\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)S_{w_{c}^{2}}^{(N_{1})}.$ (69)
@@
$\lim_{N_{1}\rightarrow\infty}\textrm{Var}_{\chi}\left({O_{c}}\right)=\sigma_{w}^{2}\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)\;\;\forall c.$ (70)
@@
  • $\left\langle{U^{(1;k)}}\right\rangle=0$:
@@
    $\lim_{N_{1}\rightarrow\infty}f_{O_{c}}^{(\chi)}\left({o}\right)=\mathcal{N}\left(o;0,\sigma_{w}^{2}\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)\right).$ (71)
  • $\left\langle{U^{(1;k)}}\right\rangle\neq 0$:
@@
    $\lim_{N_{1}\rightarrow\infty}f_{O_{c}}^{(\chi)}\left({o}\right)=\mathcal{N}\left(o;\left\langle{U^{(1;k)}}\right\rangle S_{w_{c}}^{(N_{1})},\sigma_{w}^{2}\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)\right).$ (72)
@@
    $f_{\mu_{c}}\left({m}\right)=\mathcal{N}\left(m;0,\sigma_{w}^{2}\left\langle{U^{(1;m)}}\right\rangle^{2}\right).$ (74)
D.2 Derivation of $\boldsymbol{f_{\mu_{c}}\left({m}\right)}$ for some emblematic cases
Thm. D.1 demonstrates that the decisive element for deriving $f_{\mu_{c}}\left({m}\right)$ is $\left\langle{U^{(1;k)}}\right\rangle$. To make the discussion more concrete, we will consider some cases to illustrate their distinguishing features. In particular, we will use the same settings discussed at the beginning of Sec. 5.
@@ D.2 Derivation of $\boldsymbol{f_{\mu_{c}}\left({m}\right)}$ for some emblematic cases
$U^{(1;1)}_{i}=U^{(1;1)}\left({H\left({Z^{(1)}_{i}}\right)}\right)=Z^{(1)}_{i}$. (75)
From Eq.s (62) and (75), it follows that $\left\langle{U^{(1;1)}}\right\rangle=0$. Therefore, in this setting, we have no presence of IGB, and according to Eq. 60, we have:
$\lim_{D\rightarrow\infty}f_{\mu_{c}}\left({m}\right)=\delta(m).$ (76)
Note that (76) is formally true only in the limit of an infinite size dataset. For a finite size dataset, we must compute the average $\left\langle{U^{(1;1)}}\right\rangle$ using the empirical distribution, $f_{Z^{(1)}_{k}}^{(\chi;E)}\left({z{}{}}\right)\equiv\sum_{a=1}^{D}\frac{1}{D}\delta\left(z{}{}-\hat{h}_{k}^{(1)}\left(\boldsymbol{\xi}^{(a)}\right)\right)$, where the set $\left\{\hat{h}_{k}^{(1)}\left(\boldsymbol{\xi}^{(a)}\right)\right\}_{a=1}^{D}$ contains the values mapped from each element of the dataset to the $k^{\text{th}}$ node of the first hidden layer. We then have
$\begin{split}&\left\langle{U^{(1;1)}_{j}}\right\rangle=\int_{\mathbb{R}}z{}{}\,f_{Z^{(1)}_{k}}^{(\chi;E)}\left({z{}{}}\right)dz{}{}=\frac{1}{D}\sum_{a=1}^{D}\hat{h}_{j}^{(1)}\left(\boldsymbol{\xi}^{(a)}\right)\underset{\boldsymbol{A}}{\simeq}\frac{1}{\sqrt{D}}Z_{j}=\tilde{Z}_{j},\end{split}$ (77)
@@
Therefore, considering finite size effects, $\left\langle{U^{(1;1)}}\right\rangle\neq\left\langle{U^{(1;1)}_{j}}\right\rangle$. In this case, instead of starting from Eq. (68), we can proceed similarly to the derivation of Eq. (LABEL:eq:<h>^l+1), obtaining
$f_{\mu_{c}}\left({m}\right)=\mathcal{N}\left(m;0,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\tilde{Z}_{j}^{2}\mkern-1.5mu}\mkern 1.5mu\right)=\mathcal{N}\left(x;0,\frac{4}{D}\right).$ (78)
Note that in step $\boldsymbol{A}$ of Eq. (77), we again used the CLT for the sum of r.v.s. The key difference with the previous derivation is that we do not use the CLT for $f_{Z^{(1)}_{k}}^{(\chi)}\left({z{}{}}\right)$; instead, we substitute the empirical distribution, $f_{Z^{(1)}_{k}}^{(\chi;E)}\left({z{}{}}\right)$. This latter distribution, defined on the dataset elements, accounts for the finite size effects of the dataset itself (see dependence from $D$). Note that the distribution of the variables involved in the sum in the last part of Eq. (77) ($\tilde{Z}_{j}$) narrows as $D\rightarrow\infty$; in this limit, the finite size effects disappear, and we converge to the result in Eq. (76).
@@
$U^{(1;1)}_{i}=U^{(1;1)}\left({H^{(1)}_{j}}\right)=H^{(1)}_{j}=H\left({Z^{(1)}_{j}}\right)\equiv\max\{0,Z^{(1)}_{j}\}$. (79)
$H^{(1)}_{j}$ will follow the same distribution as $Z^{(1)}_{j}$ on the positive support, since $H^{(1)}_{j}=Z^{(1)}_{j}$ for $Z^{(1)}_{j}>0$. The probability density on the negative support of $f_{Z^{(1)}_{j}}^{(\chi)}\left({z{}{}}\right)$ will collapse to 0 since $H^{(1)}_{j}=0$ for $Z^{(1)}_{j}<0$. Substituting $f_{Z^{(1)}_{j}}^{(\chi)}\left({z{}{}}\right)$ with the Gaussian distribution from Eq. (63), we get:
$\begin{split}&f_{H^{(1)}_{j}}^{(\chi)}\left({h{}{}}\right)=\left(\int_{\mathbb{R}^{-}}f_{Z^{(1)}_{j}}^{(\chi)}\left({z{}{}}\right)dz{}{}\right)\delta\left(h{}{}\right)+\Theta\left(h{}{}\right)f_{Z^{(1)}_{j}}^{(\chi)}\left({z{}{}}\right)=\frac{1}{2}\delta\left(z{}{}\right)+\Theta\left(z{}{}\right)\mathcal{N}\left(z{}{};0,\sigma_{w}^{2}\right),\end{split}$ (80)